OpenStudy (anonymous):

Solve the system of equations. 5x + y = 13 3x = 15 – 3y

OpenStudy (anonymous):

use substitution method

OpenStudy (anonymous):

5x + y = 13..............(i) 3x + 3y = 15.............(ii) therefore: from (i) y= 13-5x substituting in (ii) 3x + 3(13-5x) =15 3x +39 - 15x = 15 -12x = -24 dividing all through by -12 x = 2 and y = 13 - 5x y = 13 - 5(2) y = 13 -10 y = 3

OpenStudy (anonymous):

thank youu

OpenStudy (anonymous):

welcome